OpenStudy (anonymous):

A pallet requires 2 minutes and 15 seconds to prepare for shipment, and 30 seconds to load on a trailer. How long will it take to prepare 40 pallets and load them onto a trailer

OpenStudy (haseeb96):

just multiply 40 into 30 seconds =1200 secs require to load them onto a trailer ... if u got it then click on best respoinse

OpenStudy (anonymous):

@haseeb96 the question is "how long will it take to prepare and load 40 pallets" not just loading..

OpenStudy (anonymous):

so for each pallet to prepare and load it takes 165 seconds to prepare=2min 15 sec=2*60 sec+15=120+15=135 to load=30 sec so total to prepare and load a pellet it takes 135+30 sec=165 sec

OpenStudy (anonymous):

now calculate it for 40 pallets 40*165sec=6600 sec to convert into minutes divide it by 60 6600/60=110 minutes=1hr50 minutes

OpenStudy (anonymous):

so it takes 1hr 50 minutes to prepare and load 40 pallets